OpenStudy (anonymous):

(5x+3)25x^2-9/5x^2+37x-24 divide

OpenStudy (mertsj):

Should there be parentheses around 25x^2-9 ?

OpenStudy (mertsj):

\[\frac{(5x+3)(25x^2-9)}{5x^2+37x-24}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

simplify your answer

OpenStudy (mertsj):

What is the problem then?

OpenStudy (anonymous):

reverse the problem

OpenStudy (mertsj):

Because this is what you have posted: \[(5x+3)25x^2-\frac{9}{5x^2}+37x-24\]

OpenStudy (mertsj):

That is the problem you have posted. Do you care to correct it?

OpenStudy (anonymous):

i guess you put the numerator on the bottom and the denominator on top

OpenStudy (mertsj):

Well, no one can help you until you post the problem with which you want help.

OpenStudy (anonymous):

then you simplify your answer

OpenStudy (mertsj):

Post the correct problem

OpenStudy (anonymous):

i did you have to reverse the problem to where the numerator on bottom and denominator is on top

OpenStudy (mertsj):

Ok. Her goes my last guess as to what problem you want help with:

OpenStudy (anonymous):

thank you for all your help

OpenStudy (mertsj):

\[\frac{5x^2+37x-24}{(5x+3)(25x^2-9)}\]

OpenStudy (mertsj):

Is that the problem or not?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Can you factor?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Where did you get this problem and why are you expected to simplify it if you have not been taught to factor?

OpenStudy (anonymous):

I'm in a distant learning kind of thing taking college classes

OpenStudy (anonymous):

that is why i am on open study right now

OpenStudy (whpalmer4):

It's great that distant learning allows you to take those courses, but you're going to suffer if you don't have the necessary background! You might consider watching these videos to plug that hole: https://www.khanacademy.org/search?page_search_query=factoring

OpenStudy (mertsj):

To factor the numerator, multiply the 5 and the -24 and then find 2 factors of that number whose sum is -37.

OpenStudy (anonymous):

they do not work

OpenStudy (mertsj):

What does not work?

OpenStudy (anonymous):

the web site u just gave me

OpenStudy (mertsj):

Oh. I see someone else is helping you now. Good luck.

OpenStudy (whpalmer4):

@mertsj No, keep working with her, I have to go, just pointing out a source for info on factoring

OpenStudy (anonymous):

she left

OpenStudy (whpalmer4):

Well, it's unfortunate that those videos don't play on your computer, because I have to leave as well.

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

But I'll walk you through factoring the numerator: \[5x^2+37x−24\]To factor this, multiply the first coefficient with the last one: \[5*-24 = -120\]Now you need to find two factors of -120 that add to 37. -1*120 -2*60 -3*40 hey, there's such a pair: -3 + 40 = 37 Now rewrite the middle term as two terms, using those factors as the coefficients: \[5x^2 +40x - 3x -24\]Group them with parentheses\[(5x^2+40x) -(3x+24)\]Note that the -24 changed to +24 when I moved the - outside the parentheses. Now, can you factor each of those groups?

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

what do you get?

OpenStudy (anonymous):

do not know

OpenStudy (whpalmer4):

Let's do the left one first: \[5x^2+40x\]Are there any factors common to both of those terms? If so, what are they?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

:-( Can you multiply 5x(x+8) for me?

OpenStudy (anonymous):

no don't know how

OpenStudy (whpalmer4):

Okay. This is very basic, beginning algebra. If you don't know this, there's virtually no possibility of success in the class you're seeking help for.

OpenStudy (anonymous):

just help me please

OpenStudy (whpalmer4):

Distributive property: \[a(b+c) = a*b + a*c\] \[5x(x+8) = 5x*x + 5x*8\]Can you simplify the part on the right?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

What is your math background?

OpenStudy (whpalmer4):

Well, I'll finish factoring that numerator for you: \[(5x^2+40x) -(3x+24)\]\[5x(x+8)-3(x+8)\]Now we "undo" the distributive property to give us\[(x+8)(5x-3)\] Our whole fraction is now \[\frac{(x+8)(5x-3)}{(5x+3)(25x^2−9)}\]We can factor the denominator: \[(5x+3)(25x^2-9) = (5x+3)(5x-3)(5x+3)\]because the product term \((25x^2-9)\) is a difference of squares: \[a^2-b^2 = (a+b)(a-b) = a(a-b) + b(a-b) = a*a - a*b + b*a - b*b \]\[= a^2-b^2\] Here we have \[25x^2-9\]\[a^2-b^2\]For those to match up, \[25x^2=a^2\]\[9=b^2\]so \[\sqrt{25x^2} = \sqrt{a^2}\]or\[5x=a\]and \[\sqrt{9}=\sqrt{b}\]\[3=b\]Which gives us our factoring of \[25x^2-9=a^2-b^2=(a+b)(a-b) = (5x+3)(5x-3)\] Our fraction is now\[\frac{(x+8)(5x-3)}{(5x+3)(5x+3)(5x-3)}=\frac{(x+8)\cancel{(5x-3)}}{(5x+3)(5x+3)\cancel{(5x-3)}} = \frac{x+8 } {(5x+3)^2 } \]